# EMNLP: Educator-role Moral and Normative Large Language Models Profiling

### v1

@@ Appendix E.1 RQ1 personality traits of teacher SP LLMs at All Temperatures
Model Temp ISD PP EH LD FO CA RI IQ CE CC HC PO AO Overall Reasoning-Oriented Models Claude-3.7 0.0 5.42 5.33 4.00 2.67 4.83 4.83 4.67 2.50 5.00 4.33 4.33 5.00 4.00 4.38 Claude-3.7 0.25 5.42 5.33 3.83 2.67 4.92 4.75 4.67 2.67 5.00 4.42 4.33 4.83 4.50 4.41 Claude-3.7 0.5 5.33 5.33 4.17 2.67 5.00 4.75 4.67 2.83 5.08 4.25 4.33 5.00 4.00 4.42 Claude-3.7 0.75 5.33 5.33 4.17 2.67 4.83 4.83 4.67 2.67 5.00 4.33 4.17 5.00 4.17 4.40 Claude-3.7 1.0 5.50 5.33 4.08 2.67 4.83 4.83 4.67 2.83 5.00 4.33 4.25 4.83 4.17 4.41 DeepSeek-R1 0.0 4.83 5.00 3.75 4.33 5.17 5.17 4.50 3.67 5.17 5.08 4.92 5.00 4.33 4.69 DeepSeek-R1 0.25 4.92 5.25 4.25 4.33 5.17 5.17 4.67 3.25 4.92 5.08 4.83 5.33 4.83 4.77 DeepSeek-R1 0.5 4.83 5.00 4.42 4.25 5.33 5.25 4.67 3.17 5.17 5.00 4.83 5.00 4.67 4.74 DeepSeek-R1 0.75 4.75 5.00 4.33 4.33 5.25 5.00 4.25 3.83 5.00 5.00 5.00 5.33 4.83 4.76 DeepSeek-R1 1.0 5.00 5.00 4.50 4.25 5.17 5.08 4.33 3.67 5.17 5.17 5.00 5.08 4.83 4.79 GLM-Z1-9B 0.0 6.00 5.67 2.67 2.50 5.83 5.33 4.83 3.83 5.50 3.83 5.50 5.83 5.00 4.79 GLM-Z1-9B 0.25 6.00 5.67 2.67 2.50 6.00 5.33 4.83 3.83 5.50 3.83 5.50 6.00 5.00 4.82 GLM-Z1-9B 0.5 6.00 5.67 2.67 2.67 6.00 5.83 4.83 3.83 5.42 3.92 5.50 6.00 5.00 4.87 GLM-Z1-9B 0.75 6.00 5.92 4.33 3.58 6.00 5.75 4.67 4.67 5.50 4.75 5.25 5.83 4.83 5.16 GLM-Z1-9B 1.0 6.00 5.58 4.83 3.83 6.00 5.58 4.67 4.08 5.58 5.17 5.67 5.67 4.75 5.19 QwQ-32B 0.0 5.67 5.67 4.00 4.17 6.00 5.67 4.33 3.58 5.67 5.17 5.33 5.83 4.67 5.06 QwQ-32B 0.25 5.67 5.92 4.50 4.33 6.00 5.67 4.25 3.83 5.50 4.58 5.33 6.00 3.92 5.04 QwQ-32B 0.5 5.50 5.83 4.25 4.33 5.83 5.67 4.33 4.00 5.17 4.58 5.33 6.00 4.58 5.03 QwQ-32B 0.75 5.50 6.00 4.50 4.33 6.00 5.58 4.33 4.00 5.50 5.08 5.17 6.00 4.00 5.08 QwQ-32B 1.0 5.17 5.83 4.33 4.17 6.00 5.50 4.50 4.17 5.33 5.33 5.33 6.00 4.42 5.08 R1-Distill-70B 0.0 5.17 5.25 4.58 4.50 5.25 5.00 4.33 3.83 5.25 5.00 4.83 5.00 5.00 4.85 R1-Distill-70B 0.25 5.25 5.17 4.83 4.58 5.17 5.25 4.58 4.00 5.00 5.00 4.67 5.00 5.08 4.89 R1-Distill-70B 0.5 5.33 5.17 4.83 4.58 5.00 5.00 4.67 3.83 5.08 5.00 4.83 5.08 4.92 4.87 R1-Distill-70B 0.75 5.17 5.25 4.75 4.67 5.00 5.00 4.67 4.17 5.00 5.00 4.83 5.33 5.08 4.92 R1-Distill-70B 1.0 5.33 5.25 4.83 4.50 5.00 5.00 4.67 4.00 5.00 5.00 4.83 5.33 5.00 4.90 R1-Distill-8B 0.0 5.33 5.67 5.00 4.92 5.17 5.25 4.83 4.33 4.83 5.00 4.58 4.83 4.92 4.97 R1-Distill-8B 0.25 5.00 5.33 4.75 4.83 5.00 5.17 4.83 4.67 5.08 5.00 4.83 5.17 5.50 5.01 R1-Distill-8B 0.5 5.33 5.17 4.58 4.50 4.92 5.00 4.67 4.67 5.00 4.83 4.83 5.00 4.92 4.88 R1-Distill-8B 0.75 4.92 5.33 5.00 4.58 5.08 5.17 4.83 4.50 5.17 4.92 4.58 4.83 4.83 4.90 R1-Distill-8B 1.0 5.08 5.58 4.83 4.67 5.00 5.17 4.92 4.50 5.33 4.67 4.75 4.83 4.83 4.94 Non-Reasoning Models Baichuan2-7B 0.0 4.67 3.17 4.33 3.83 3.33 3.83 3.67 2.83 2.50 3.83 3.67 3.50 4.17 3.64 Baichuan2-7B 0.25 4.33 3.67 4.58 4.00 2.92 3.83 3.00 3.83 2.83 4.50 3.25 3.50 4.33 3.74 Baichuan2-7B 0.5 4.42 3.75 3.42 4.33 3.08 3.83 3.33 4.00 4.25 3.92 3.50 3.58 4.75 3.86 Baichuan2-7B 0.75 4.17 3.58 4.75 3.83 2.17 4.17 3.67 4.00 4.25 4.67 4.58 4.17 3.17 3.94 Baichuan2-7B 1.0 4.83 3.42 4.17 3.50 4.17 4.92 4.08 3.83 3.92 4.17 5.17 3.25 5.00 4.19 DeepSeek-V3 0.0 5.17 5.33 4.83 4.33 5.00 5.00 4.83 3.67 5.00 5.33 5.17 5.00 5.17 4.91 DeepSeek-V3 0.25 5.17 5.25 4.83 4.33 5.00 5.00 4.83 3.67 5.17 5.33 5.17 5.00 5.08 4.91 DeepSeek-V3 0.5 5.17 5.33 4.83 4.33 5.00 5.00 4.83 3.67 5.17 5.33 5.17 5.00 5.08 4.92 DeepSeek-V3 0.75 5.17 5.33 4.83 4.33 5.00 5.00 4.83 3.83 5.17 5.33 5.17 5.00 5.00 4.92 DeepSeek-V3 1.0 5.17 5.33 4.83 4.33 5.00 5.00 4.83 3.58 5.08 5.33 5.17 5.00 5.08 4.90 GPT-4.1 0.0 6.00 6.00 4.33 4.00 5.83 5.92 5.08 3.83 5.83 5.17 5.50 6.00 5.75 5.33 GPT-4.1 0.25 6.00 6.00 4.50 4.00 5.83 6.00 5.17 3.83 5.83 5.33 5.50 6.00 5.83 5.37 GPT-4.1 0.5 6.00 6.00 4.50 4.00 5.83 6.00 5.17 3.83 5.83 5.33 5.50 6.00 5.83 5.37 GPT-4.1 0.75 6.00 6.00 4.50 4.00 5.83 6.00 5.17 3.83 5.83 5.33 5.50 6.00 5.83 5.37 GPT-4.1 1.0 6.00 6.00 4.50 4.00 5.83 6.00 5.08 3.83 5.83 5.33 5.50 6.00 5.83 5.37 Qwen2.5-32B 0.0 4.67 5.17 3.33 3.00 4.42 3.67 3.58 3.67 4.33 3.50 4.50 4.17 4.17 4.01 Qwen2.5-32B 0.25 4.67 5.00 3.33 3.00 4.33 3.67 3.58 3.67 4.17 3.50 4.33 4.33 4.17 3.98 Qwen2.5-32B 0.5 4.67 5.00 3.33 3.00 4.50 3.67 3.50 3.67 4.17 3.67 4.42 4.33 4.17 4.01 Qwen2.5-32B 0.75 4.67 5.25 3.42 3.00 4.25 3.67 3.50 3.67 4.25 3.50 4.33 4.17 4.17 3.99 Qwen2.5-32B 1.0 4.67 5.25 3.33 3.00 4.17 3.67 3.67 3.67 4.25 3.42 4.33 4.00 4.17 3.97 Qwen2.5-72B 0.0 5.00 5.00 4.00 3.50 5.00 4.83 4.33 3.50 4.83 5.00 4.33 4.83 4.67 4.53 Qwen2.5-72B 0.25 5.00 5.00 4.00 3.50 5.00 4.83 4.33 3.67 4.83 5.00 4.33 4.83 4.67 4.54 Qwen2.5-72B 0.5 5.00 5.00 4.00 3.50 5.00 4.83 4.33 3.58 4.75 5.00 4.33 4.83 4.67 4.53 Qwen2.5-72B 0.75 5.00 5.00 4.00 3.50 5.00 4.83 4.33 3.67 4.83 5.00 4.33 4.83 4.67 4.54 Qwen2.5-72B 1.0 5.00 5.00 4.00 3.50 5.00 4.83 4.33 3.50 4.67 5.00 4.33 4.83 4.67 4.51 Qwen2.5-7B 0.0 5.00 5.00 4.83 3.58 5.00 5.00 4.67 2.83 5.00 5.00 5.00 4.67 4.67 4.63 Qwen2.5-7B 0.25 5.00 5.00 4.83 3.50 5.00 5.00 4.67 2.83 5.00 5.00 5.00 4.67 4.83 4.64 Qwen2.5-7B 0.5 5.00 5.00 4.67 3.50 5.00 5.00 4.67 3.17 5.00 5.00 5.00 4.67 4.83 4.65 Qwen2.5-7B 0.75 5.00 5.00 4.83 3.58 5.00 5.00 4.67 3.17 5.00 5.00 5.00 4.75 4.83 4.68 Qwen2.5-7B 1.0 5.00 5.00 4.83 3.67 5.00 5.00 4.67 3.25 5.00 5.00 5.00 4.67 4.67 4.67
Appendix E.2 RQ2 LLM Stages of Moral Development at All Temperatures
@@
The results in Table 11 show how models’ moral development stage distributions vary with temperature, revealing patterns in the stability and variability of teacher-role LLMs’ moral reasoning performance across different generation configurations.
@@
Table 11: MSS of LLMs Across 5 Moral Dimensions (temperature $t\!=\!0\text{\textendash}1$) with Heat-map Coloring and Overall Average (88 dilemmas)

### v2

@@ Appendix H.1 RQ1 personality traits of teacher SP LLMs at All Temperatures
Model Temp ISD PP EH LD FO CA RI IQ CE CC HC PO AO Overall Reasoning-Oriented Models Claude-3.7 0.0 5.42 5.33 4.00 2.67 4.83 4.83 4.67 2.50 5.00 4.33 4.33 5.00 4.00 4.38 0.25 5.42 5.33 3.83 2.67 4.92 4.75 4.67 2.67 5.00 4.42 4.33 4.83 4.50 4.41 0.5 5.33 5.33 4.17 2.67 5.00 4.75 4.67 2.83 5.08 4.25 4.33 5.00 4.00 4.42 0.75 5.33 5.33 4.17 2.67 4.83 4.83 4.67 2.67 5.00 4.33 4.17 5.00 4.17 4.40 1.0 5.50 5.33 4.08 2.67 4.83 4.83 4.67 2.83 5.00 4.33 4.25 4.83 4.17 4.41 DeepSeek-R1 0.0 4.83 5.00 3.75 4.33 5.17 5.17 4.50 3.67 5.17 5.08 4.92 5.00 4.33 4.69 0.25 4.92 5.25 4.25 4.33 5.17 5.17 4.67 3.25 4.92 5.08 4.83 5.33 4.83 4.77 0.5 4.83 5.00 4.42 4.25 5.33 5.25 4.67 3.17 5.17 5.00 4.83 5.00 4.67 4.74 0.75 4.75 5.00 4.33 4.33 5.25 5.00 4.25 3.83 5.00 5.00 5.00 5.33 4.83 4.76 1.0 5.00 5.00 4.50 4.25 5.17 5.08 4.33 3.67 5.17 5.17 5.00 5.08 4.83 4.79 GLM-Z1-9B 0.0 6.00 5.67 2.67 2.50 5.83 5.33 4.83 3.83 5.50 3.83 5.50 5.83 5.00 4.79 0.25 6.00 5.67 2.67 2.50 6.00 5.33 4.83 3.83 5.50 3.83 5.50 6.00 5.00 4.82 0.5 6.00 5.67 2.67 2.67 6.00 5.83 4.83 3.83 5.42 3.92 5.50 6.00 5.00 4.87 0.75 6.00 5.92 4.33 3.58 6.00 5.75 4.67 4.67 5.50 4.75 5.25 5.83 4.83 5.16 1.0 6.00 5.58 4.83 3.83 6.00 5.58 4.67 4.08 5.58 5.17 5.67 5.67 4.75 5.19 QwQ-32B 0.0 5.67 5.67 4.00 4.17 6.00 5.67 4.33 3.58 5.67 5.17 5.33 5.83 4.67 5.06 0.25 5.67 5.92 4.50 4.33 6.00 5.67 4.25 3.83 5.50 4.58 5.33 6.00 3.92 5.04 0.5 5.50 5.83 4.25 4.33 5.83 5.67 4.33 4.00 5.17 4.58 5.33 6.00 4.58 5.03 0.75 5.50 6.00 4.50 4.33 6.00 5.58 4.33 4.00 5.50 5.08 5.17 6.00 4.00 5.08 1.0 5.17 5.83 4.33 4.17 6.00 5.50 4.50 4.17 5.33 5.33 5.33 6.00 4.42 5.08 R1-Distill-70B 0.0 5.17 5.25 4.58 4.50 5.25 5.00 4.33 3.83 5.25 5.00 4.83 5.00 5.00 4.85 0.25 5.25 5.17 4.83 4.58 5.17 5.25 4.58 4.00 5.00 5.00 4.67 5.00 5.08 4.89 0.5 5.33 5.17 4.83 4.58 5.00 5.00 4.67 3.83 5.08 5.00 4.83 5.08 4.92 4.87 0.75 5.17 5.25 4.75 4.67 5.00 5.00 4.67 4.17 5.00 5.00 4.83 5.33 5.08 4.92 1.0 5.33 5.25 4.83 4.50 5.00 5.00 4.67 4.00 5.00 5.00 4.83 5.33 5.00 4.90 R1-Distill-8B 0.0 5.33 5.67 5.00 4.92 5.17 5.25 4.83 4.33 4.83 5.00 4.58 4.83 4.92 4.97 0.25 5.00 5.33 4.75 4.83 5.00 5.17 4.83 4.67 5.08 5.00 4.83 5.17 5.50 5.01 0.5 5.33 5.17 4.58 4.50 4.92 5.00 4.67 4.67 5.00 4.83 4.83 5.00 4.92 4.88 0.75 4.92 5.33 5.00 4.58 5.08 5.17 4.83 4.50 5.17 4.92 4.58 4.83 4.83 4.90 1.0 5.08 5.58 4.83 4.67 5.00 5.17 4.92 4.50 5.33 4.67 4.75 4.83 4.83 4.94 Qwen3-235B-R 0.0 5.33 5.00 5.00 5.00 5.00 5.00 5.00 4.17 5.00 5.00 5.17 5.00 5.00 4.97 0.25 5.33 5.17 5.00 5.00 5.00 5.00 5.00 4.17 5.00 5.00 5.17 5.00 5.00 4.99 0.5 5.33 5.08 5.00 5.00 5.00 5.00 5.00 4.17 5.00 5.00 5.17 5.00 5.00 4.98 0.75 5.33 5.17 5.00 5.00 5.00 5.00 5.00 4.17 5.00 5.00 5.17 5.00 5.00 4.99 1.0 5.25 5.00 5.00 5.00 5.00 5.00 5.00 4.17 5.00 5.00 5.00 5.00 5.00 4.96 Non-Reasoning Models Baichuan2-7B 0.0 4.67 3.17 4.33 3.83 3.33 3.83 3.67 2.83 2.50 3.83 3.67 3.50 4.17 3.64 0.25 4.33 3.67 4.58 4.00 2.92 3.83 3.00 3.83 2.83 4.50 3.25 3.50 4.33 3.74 0.5 4.42 3.75 3.42 4.33 3.08 3.83 3.33 4.00 4.25 3.92 3.50 3.58 4.75 3.86 0.75 4.17 3.58 4.75 3.83 2.17 4.17 3.67 4.00 4.25 4.67 4.58 4.17 3.17 3.94 1.0 4.83 3.42 4.17 3.50 4.17 4.92 4.08 3.83 3.92 4.17 5.17 3.25 5.00 4.19 DeepSeek-V3 0.0 5.17 5.33 4.83 4.33 5.00 5.00 4.83 3.67 5.00 5.33 5.17 5.00 5.17 4.91 0.25 5.17 5.25 4.83 4.33 5.00 5.00 4.83 3.67 5.17 5.33 5.17 5.00 5.08 4.91 0.5 5.17 5.33 4.83 4.33 5.00 5.00 4.83 3.67 5.17 5.33 5.17 5.00 5.08 4.92 0.75 5.17 5.33 4.83 4.33 5.00 5.00 4.83 3.83 5.17 5.33 5.17 5.00 5.00 4.92 1.0 5.17 5.33 4.83 4.33 5.00 5.00 4.83 3.58 5.08 5.33 5.17 5.00 5.08 4.90 GPT-4.1 0.0 6.00 6.00 4.33 4.00 5.83 5.92 5.08 3.83 5.83 5.17 5.50 6.00 5.75 5.33 0.25 6.00 6.00 4.50 4.00 5.83 6.00 5.17 3.83 5.83 5.33 5.50 6.00 5.83 5.37 0.5 6.00 6.00 4.50 4.00 5.83 6.00 5.17 3.83 5.83 5.33 5.50 6.00 5.83 5.37 0.75 6.00 6.00 4.50 4.00 5.83 6.00 5.17 3.83 5.83 5.33 5.50 6.00 5.83 5.37 1.0 6.00 6.00 4.50 4.00 5.83 6.00 5.08 3.83 5.83 5.33 5.50 6.00 5.83 5.37 Qwen2.5-32B 0.0 4.67 5.17 3.33 3.00 4.42 3.67 3.58 3.67 4.33 3.50 4.50 4.17 4.17 4.01 0.25 4.67 5.00 3.33 3.00 4.33 3.67 3.58 3.67 4.17 3.50 4.33 4.33 4.17 3.98 0.5 4.67 5.00 3.33 3.00 4.50 3.67 3.50 3.67 4.17 3.67 4.42 4.33 4.17 4.01 0.75 4.67 5.25 3.42 3.00 4.25 3.67 3.50 3.67 4.25 3.50 4.33 4.17 4.17 3.99 1.0 4.67 5.25 3.33 3.00 4.17 3.67 3.67 3.67 4.25 3.42 4.33 4.00 4.17 3.97 Qwen2.5-72B 0.0 5.00 5.00 4.00 3.50 5.00 4.83 4.33 3.50 4.83 5.00 4.33 4.83 4.67 4.53 0.25 5.00 5.00 4.00 3.50 5.00 4.83 4.33 3.67 4.83 5.00 4.33 4.83 4.67 4.54 0.5 5.00 5.00 4.00 3.50 5.00 4.83 4.33 3.58 4.75 5.00 4.33 4.83 4.67 4.53 0.75 5.00 5.00 4.00 3.50 5.00 4.83 4.33 3.67 4.83 5.00 4.33 4.83 4.67 4.54 1.0 5.00 5.00 4.00 3.50 5.00 4.83 4.33 3.50 4.67 5.00 4.33 4.83 4.67 4.51 Qwen2.5-7B 0.0 5.00 5.00 4.83 3.58 5.00 5.00 4.67 2.83 5.00 5.00 5.00 4.67 4.67 4.63 0.25 5.00 5.00 4.83 3.50 5.00 5.00 4.67 2.83 5.00 5.00 5.00 4.67 4.83 4.64 0.5 5.00 5.00 4.67 3.50 5.00 5.00 4.67 3.17 5.00 5.00 5.00 4.67 4.83 4.65 0.75 5.00 5.00 4.83 3.58 5.00 5.00 4.67 3.17 5.00 5.00 5.00 4.75 4.83 4.68 1.0 5.00 5.00 4.83 3.67 5.00 5.00 4.67 3.25 5.00 5.00 5.00 4.67 4.67 4.67 Qwen3-235B-C 0.0 5.33 5.00 5.00 5.00 5.00 5.00 5.00 4.17 5.00 5.00 5.17 5.00 5.00 4.97 0.25 5.33 5.08 5.00 5.00 5.00 5.00 5.00 4.17 5.00 5.00 5.17 5.00 5.00 4.98 0.5 5.33 5.00 5.00 5.00 5.00 5.00 5.00 4.17 5.00 5.00 5.08 5.00 5.00 4.97 0.75 5.33 5.17 5.00 5.00 5.00 5.00 5.00 3.92 5.00 5.00 5.17 5.00 5.00 4.97 1.0 5.33 5.00 5.00 5.00 5.00 5.00 5.00 3.92 5.00 5.00 5.17 5.00 5.00 4.96
Appendix H.2 RQ2 LLM Stages of Moral Development at All Temperatures
@@
The results in Table 15 and  16show how models’ moral development stage distributions vary with temperature, revealing patterns in the stability and variability of teacher-role LLMs’ moral reasoning performance across different generation configurations.
@@
Table 15: MSS of 14 LLMs Across 5 Moral Dimensions (English, temperature $t\!=\!0\text{\textendash}1$) with Heat-map Coloring and Overall Average (88 dilemmas)